# Avoiding Catastrophe in Continuous Spaces by Asking for Help

### v2

@@ B.1 Proof roadmap
    For each $t\in[T]$, let $X_{t},$ refer to the values of $X$ respectively at the start of time step $t$. Define $\pi_{t}$ similarly, except only for $t>g(T)$.
@@
    Let $V_{T}=\{t\in[T]:\mu(s_{t},a_{t})<\mu^{*}(s_{t})\}$ be the set of time steps with a suboptimal payoff.
@@
    Unless specified (e.g., $\Pr_{s\sim\mathcal{D}}$), all probabilities are with respect to $\mathbf{s}\sim\mathcal{D}^{T}$.
The proof proceeds in the following steps:
  1. 1.
    Show that $\pi_{t}$ meets the criteria for the fundamental theorem of multi-class PAC learning and thus has small expected error on a single state (Lemma B.1).
  2. 2.
    Use Lemma B.1 and a non-standard concentration inequality (Lemma B.2) to show that $|V_{T}|$ is small with high probability. This involves two steps (Lemmas B.3 and B.4).
  3. 3.
    Having a small number of mistakes is not sufficient, because a single payoff of $\mu(s_{t},a_{t})=0$ can ruin the entire endeavor. Lemmas B.5 and B.6 bound the suboptimality of a single payoff.
@@
    Lemma B.7 transforms the multiplicative objective into an additive form that is easier to manipulate.
@@
    Lemma B.8 bounds the total regret of the algorithm.
  6. 6.
    Lemma B.10 bounds the total number of queries based on a packing number lemma (Lemma B.9).
Lemmas B.8 and B.10 together imply Theorem 4.5.
B.2 Proof
@@
With probability at least $1-1/T$, we have $\Pr_{s\sim\mathcal{D}}[\pi_{t}(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$ $\forall t>g(T)$.
@@
Let $\delta=1/T$. Then
$|X_{g(T)+1}|=$ $\ g(T)$
$=$ $\ \frac{cd}{\varepsilon(T)}\log(d\,T)$
$=$ $\ 2\frac{cd}{\varepsilon(T)}\log(d^{2}T^{2})$
$=$ $\ \frac{c}{\varepsilon(T)}(d\log(d^{2}T^{2})+d\log(d^{2}T^{2}))$
$\geq$ $\ \frac{c}{\varepsilon(T)}(d\log(d\,T\cdot T^{\frac{n}{n+0.5}})+\log T)$
$\geq$ $\ \frac{c}{\varepsilon(T)}\left(d\log(\frac{d|A|}{\varepsilon(T)})+\log\frac{1}{\delta}\right)$
Now consider any $t>g(T)$. By construction, $\pi_{t}$ is $(X_{t},r)$-local ($\pi$ and $X$ are not updated on every time step, but are always updated at the same time). Therefore $\pi_{t}$ is consistent with $X_{t}$ and in turn is consistent with $X_{g(T)+1}$ since $X_{g(T)+1}\subseteq X_{t}$. Since $X_{g(T)+1}$ is a set of i.i.d. samples of the requisite size, Theorem 4.4 implies that with at probability at least $1-1/T$, any policy $\pi\in\Pi$ consistent with $X_{g(T)+1}$ satisfies $\Pr_{s\sim\mathcal{D}}[\pi(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$. In particular, this holds for $\pi_{t}$ for all $t>g(T)$, completing the proof. ∎
Next, we wish to prove that $|V_{T}|$ is small. The $\{t\in V_{T}\}$ indicator variables are not independent, and standard martingale concentration bounds like the Azuma-Hoeffding inequality are not tight enough, so we use a specialized inequality from [8]. The statement below is slightly simplified from its original source, since we do not need the more general version.
Lemma B.2 (Corollary 2.1 in [8]).
Let $x_{0},\dots,x_{k}$ be a supermartingale and assume there exists $b\in(0,1]$ such that $x_{i}-x_{i-1}\in[-b,1]$ for all $i\in[k]$. Then for any $x>0$,
$\Pr[x_{k}\geq x]\leq\exp\left(-\frac{x^{2}}{2(kb+x/3)}\right)$
Lemma B.3.
Assume $\Pr_{s\sim\mathcal{D}}[\pi_{t}(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$ for all $t>g(T)$. Then
$\Pr\left[|V_{T}|\geq 2T\varepsilon(T)\right]\leq\exp(-3T\varepsilon(T)/8)$
Proof.
Let $k=T-g(T)$ for brevity. For each $t\in\{g(T)+1,\dots,T\}$, define a random variable $y_{t-g(T)}=\mathbf{1}_{\pi_{t}(s_{t})\neq\pi^{*}(s_{t})}-\varepsilon(T)$. Also let $x_{0}=0$, and for $i\in[k]$, let $x_{i}=x_{i-1}+y_{i}$. This defines random variables $y_{1},\dots,y_{k}$ and $x_{0},\dots,x_{k}$. For each $i\in[k]$, let $y_{<i}=\{y_{1},\dots,y_{i-1}\}$ and $x_{<i}=\{x_{0},\dots,x_{i-1}\}$.
To show that $x_{0},\dots,x_{k}$ form a supermartingale, we first claim that $\operatorname*{\mathbb{E}}[y_{i}\mid y_{<i}]\leq 0$ for all $i\in[k]$. Regardless of what occurs on previous time steps, each state is sampled randomly from $\mathcal{D}$. Thus
$\operatorname*{\mathbb{E}}[y\mid y_{<i}]=\Pr_{s\sim\mathcal{D}}[\pi_{i+g(T)}\neq\pi^{*}(s)\mid y_{<i}]-\varepsilon(T)\leq 0$
Since $y_{<i}$ and $x_{<i}$ provide the same information (i.e., either can be constructed from the other), we have $\operatorname*{\mathbb{E}}[y_{i}\mid x_{<i}]\leq 0$. Therefore
$\operatorname*{\mathbb{E}}[x_{i}\mid x_{<i}]=$ $\ \operatorname*{\mathbb{E}}[x_{i-1}+y_{i}-\varepsilon(T)\mid x_{<i}]$
$=$ $\ \operatorname*{\mathbb{E}}[x_{i-1}\mid x_{<i}]+\operatorname*{\mathbb{E}}[y_{i}\mid x_{<i}]-\varepsilon(T)$
$\leq$ $\ x_{i-1}+\varepsilon(T)-\varepsilon(T)$
$=$ $\ x_{i-1}$
Also, since each $x_{i}$ is bounded, we have $\operatorname*{\mathbb{E}}[x_{i}]<\infty$ for all $i\in[k]$. Therefore $x_{0},\dots,x_{k}$ is indeed a supermartingale. Furthermore, $x_{i}-x_{i-1}=y_{i}\in[-\varepsilon(T),1]$ for all $i\in[k]$, so by Lemma B.2,
@@
$\Pr[x_{k}\geq 2T\varepsilon(T)]\leq$ $\ \exp\left(-\frac{T^{2}\varepsilon(T)^{2}}{2(k\varepsilon(T)+T\varepsilon(T)/3)}\right)$
Since $k\leq T$,
$\Pr[x_{k}\geq 2T\varepsilon(T)]\leq$ $\ \exp\left(-\frac{T^{2}\varepsilon(T)^{2}}{2(T\varepsilon(T)+T\varepsilon(T)/3)}\right)$
$=$ $\ \exp\left(-\frac{T^{2}\varepsilon(T)^{2}}{8T\varepsilon(T)/3}\right)$
$=$ $\ \exp(-3T\varepsilon(T)/8)$
@@
Let $\xi$ be the event that $\Pr_{s\sim\mathcal{D}}[\pi_{t}(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$ for all $t>g(T)$. Lemma B.1 states that $\Pr[\neg\xi]\leq 1/T$ and Lemma B.3 states that $\Pr[|V_{T}|\geq 2T\varepsilon(T)\mid\xi]\leq\exp(-3T\varepsilon(T)/8)$. Then
$\Pr[|V_{T}|\geq 2T\varepsilon(T)]=$ $\ \Pr[|V_{T}|\geq 2T\varepsilon(T)\mid\xi]\Pr[\xi]+\Pr[|V_{T}|\geq 2T\varepsilon(T)\mid\neg\xi]\Pr[\neg\xi]$
$\leq$ $\ \exp(-3T\varepsilon(T)/8)\cdot 1+1\cdot 1/T$
$=$ $\ \exp(-3T^{1-\frac{n}{n+0.5}}/8)+1/T$
$=$ $\ \exp(-3T^{\frac{0.5}{n+0.5}}/8)+1/T$
as claimed. ∎
Remark B.1.
The astute reader note that Lemma B.5 is the only place in the paper where we invoke the assumption that the mentor’s policy is optimal. Both of our positive results (Theorems 4.5 and 5.2) can be generalized to a possibly suboptimal mentor policy $\pi^{m}$ if:
  1. 1.
    We modify the Lipschitz assumption to be $|\mu(s,\pi^{m}(s))-\mu(s,\pi^{m}(s^{\prime}))|\leq L||s-s^{\prime}||$. In words, the value of the mentor’s policy shouldn’t change too much between nearby states.
  2. 2.
    We define replace $\mu^{*}(s)$ with $\mu(s,\pi^{m}(s))$ throughout.
Lemma B.5 follows immediately from these two assumptions.
Also note that Lemmas B.5 and B.7 are not specific to this algorithm and will also be used in the proof of Theorem 5.2.
Lemma B.5.
If $||s-s^{\prime}||\leq\delta$ for some $\delta\geq 0$, then $\mu(s,\pi^{*}(s^{\prime}))\geq\mu^{*}(s)-2L\delta$.
Proof.
We have By the Lipschitz continuity of $\mu$, $\mu(s,\pi^{*}(s^{\prime}))\geq\mu(s^{\prime},\pi^{*}(s^{\prime}))-L||s-s^{\prime}||$. Also, $\mu(s^{\prime},\pi^{*}(s^{\prime}))=\max_{a\in A}\mu(s^{\prime},a)\geq\mu(s^{\prime},\pi^{*}(s))$ by assumption. Thus
$\mu(s,\pi^{*}(s^{\prime}))\geq$ $\ \mu(s^{\prime},\pi^{*}(s^{\prime}))-L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$)
$\geq$ $\ \mu(s^{\prime},\pi^{*}(s))-L||s-s^{\prime}||$ (By assumption, $\pi^{*}$ is optimal)
$\geq$ $\ \mu(s,\pi^{*}(s))-2L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$ again)
$=$ $\ \mu^{*}(s)-2L||s-s^{\prime}||$ (Definition of $\mu^{*}(s)$)
Plugging in $||s-s^{\prime}||\leq\delta$ completes the proof. ∎
Lemma B.6.
Under the conditions of Theorem 4.5, $\mu(s_{t},a_{t})\geq\mu^{*}(s_{t})-2Lr$ for all $t\in[T]$.
Proof.
If we query the mentor at time $t$, then $\mu(s_{t},a_{t})=\mu^{*}(s_{t})$, so assume we do not query the mentor at time $t$. Let $(s^{\prime},a^{\prime})=\operatorname*{arg\,min}_{(s,a)\in X_{t}}||s_{t}-s||$: then $||s_{t}-s^{\prime}||\leq r$. Since $\pi_{t}$ is $(X_{t},r)$-local by construction, there exists $(s^{\prime\prime},a^{\prime\prime})\in X_{t}$ such that $||s_{t}-s^{\prime\prime}||\leq r$ and $a_{t}=\pi_{t}(s_{t})=a^{\prime\prime}$. Then by Lemma B.5, $\mu(s_{t},\pi^{*}(s^{\prime\prime}))\geq\mu^{*}(s_{t})-2Lr$. Furthermore, because $(s^{\prime\prime},a^{\prime\prime})\in X$, we have $\pi^{*}(s^{\prime\prime})=a^{\prime\prime}=a_{t}$. Therefore $\mu(s_{t},a_{t})\geq\mu^{*}(s_{t})-2Lr$. ∎
Lemma B.7.
If $\mu^{*}(s_{t})\geq\mu_{0}>0$ and $c_{t}\in\big{[}0,\mu_{0}/2\big{]}$ for all $t\in[T]$, then
$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq\prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}$
Proof.
Recall the standard limit $\lim_{\delta\to 0}\delta^{-1}\log(1+\delta)=1$. Thus for all $\varepsilon>0$, there exists $\delta_{0}>0$ such that whenever $|\delta|\leq\delta_{0}$, we have $|\delta^{-1}\log(1+\delta)-1|\leq\varepsilon$. For our purposes, $\varepsilon=0.5$ is sufficient, we only care about $\delta<0$, and we only need the lower bound. One can verify that $\delta_{0}=0.5$ is sufficient. To summarize, we have $1.5\delta\leq\log(1+\delta)$ whenever $|\delta|\leq 0.5$.
Let $\delta_{t}=-c_{t}/\mu^{*}(s_{t})$ for each $t\in[T]$. Then
$|\delta_{t}|=\frac{c_{t}}{\mu^{*}(s_{t})}\leq\frac{\mu_{0}/2}{\mu_{0}}=0.5$
and so
$-\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \log\left(1-\frac{c_{t}}{\mu^{*}(s_{t})}\right)$
$\log(\mu^{*}(s_{t}))-\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \log\left(\mu^{*}(s_{t})-c_{t}\right)$
Now summing over $t$,
$\sum_{t=1}^{T}\log(\mu^{*}(s_{t}))-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \sum_{t=1}^{T}\log\left(\mu^{*}(s_{t})-c_{t}\right)$
$\log\prod_{t=1}^{T}\mu^{*}(s_{t})-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \log\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)$
Exponentiating both sides yields
$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq\exp\left(-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\right)\prod_{t=1}^{T}\mu^{*}(s_{t})$
Applying the standard inequality $e^{x}\geq 1+x$ gives us
$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq$ $\ \left(1-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\right)\prod_{t=1}^{T}\mu^{*}(s_{t})$
$=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}\frac{\prod_{i=1}^{T}\mu^{*}(s_{i})}{\mu^{*}(s_{t})}$
Since $\mu^{*}(s_{t})\in[0,1]$ for all $t\in[T]$, $\prod_{i=1}^{T}\mu^{*}(s_{i})\leq\mu^{*}(s_{t})$. Thus
$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}\frac{\mu^{*}(s_{t})}{\mu^{*}(s_{t})}$
$=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}$
@@
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]\leq\frac{6L}{T^{\frac{0.5}{n+0.5}}}+\frac{1}{T}+\exp(-3T^{\frac{0.5}{n+0.5}}/8)$
@@
Let $\mu_{0}=4Lr$. If there exists $t\in[T]$ with $\mu^{*}(s_{t})<\mu_{0}$, then
$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})<\mu_{0}-0=4L\varepsilon(T)^{1/n}=4LT^{\frac{-1}{n+0.5}}\leq 6LT^{\frac{-0.5}{n+0.5}}$
so the claim is satisfied. Thus assume $\mu^{*}(s_{t})\geq\mu_{0}$ for all $t\in[T]$. Let $c_{t}=0$ for $t\not\in V_{T}$ and $c_{t}=2Lr$ for $t\in V_{T}$. Then by Lemma B.6,
$\prod_{t=1}^{T}\mu(s_{t},a_{t})=$ $\ \prod_{t\in V_{T}}\mu(s_{t},a_{t})\prod_{t\not\in V_{T}}\mu(s_{t},a_{t})$
$\geq$ $\ \prod_{t\in V_{T}}\left(\mu^{*}(s_{t})-2Lr\right)\prod_{t\not\in V_{T}}\mu^{*}(s_{t})$
$=$ $\ \prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)$
Since $c_{t}\in[0,\mu_{0}/2]$ for all $t\in[T]$, Lemma B.7 implies that
$\prod_{t=1}^{T}\mu(s_{t},a_{t})\geq$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}$
$=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-3L|V_{T}|r$
If $|V_{T}|<2T\varepsilon(T)$, then
$R_{T}(\mathbf{s},\mu,\pi^{*})=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
$\leq$ $\ 6LT\varepsilon(T)r$
$=$ $\ 6LT\varepsilon(T)^{(n+1)/n}$
$=$ $\ 6LT^{(1+\frac{-n}{n+0.5}\cdot\frac{n+1}{n})}$
$=$ $\ 6LT^{(\frac{n+0.5}{n+0.5}-\frac{n+1}{n+0.5})}$
$=$ $\ 6LT^{\frac{-0.5}{n+0.5}}$
Then with the help of Lemma B.4 (and noting that $R_{T}(\mathbf{s},\mu,\pi^{*})\leq 1$ always),
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]\leq$ $\ \Pr[|V_{T}|\geq 2T\varepsilon(T)]\cdot 6LT^{\frac{-0.5}{n+0.5}}+\Pr[|V_{T}|\geq 2T\varepsilon(T)]\cdot 1$
$\leq$ $\ 6LT^{\frac{-0.5}{n+0.5}}+\frac{1}{T}+\exp(-3T^{\frac{0.5}{n+0.5}}/8)$
@@
Let $(V,||\cdot||)$ be a normed vector space and let $\delta>0$. Then $W\subseteq V$ is a $\delta$-packing of $V$ if for all $x,y\in W$, $||x-y||>\delta$. The $\delta$-packing number of $V$, denoted $\mathcal{M}(V,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $V$.
@@
Let $B$ be the unit ball and let |||||$\cdot$||| | ⋅ | | denote Euclidean distance. Assume $V\subset\mathbb{R}^{n}$ is convex, and $\delta B\subset V$ for some $\delta>0$. Then
$\mathcal{M}(V,||\cdot||,\delta)\leq\left(\frac{3}{\delta}\right)^{n}\frac{\operatorname*{vol}(V)}{\operatorname*{vol}(B)}$

### v3

@@ B.2 Main proof
    For each $t\in[T]$, let $X_{t}$ refer to the value of $X$ at the start of time step $t$.
@@
    Let $V_{T}=\{t\in[T]:\pi_{t}(s_{t})\neq\pi^{m}(s_{t})\}$ be the set of time steps where Hedge’s proposed action doesn’t match the mentor’s. Note that $|V_{T}|$ upper bounds the number of mistakes the algorithm makes (the number of mistakes could be smaller, since the algorithm sometimes queries instead of taking action $\pi_{t}(s_{t})$).
@@
    For $K\subseteq S$, let $\operatorname*{vol}(K)$ denote the $n$-dimensional Lebesgue measure of $K$.
@@
    With slight abuse of notation, we will use inequalities of the form $f(T)\leq g(T)+O(h(T))$ to mean that there exists a constant $C$ such that $f(T)\leq g(T)+Ch(T)$.
@@
    We will use “Case 1” to refer to finite VC dimension and $\sigma$-smooth $\mathbf{s}$ and “Case 2” to refer to finite Littlestone dimension. In Case 1, expectations are over the randomness of both $\mathbf{s}$ and the algorithm, while in Case 2, expectations are over just the randomness of the algorithm. When we need to distinction, we use $\operatorname*{\mathbb{E}}_{\mathbf{a}}$ to denote the expectation over randomness of the algorithm and $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}$ to denote the expectation over $\mathbf{s}$.
@@
Under the conditions of Theorem 5.1, Algorithm 1 satisfies
$\operatorname*{\mathbb{E}}[|V_{T}|]\in O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
@@
Define $\ell:[T]\times\tilde{\Pi}\to[0,1]$ by $\ell(t,\pi)=\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$, and let $w^{h}$ and $\pi_{t}^{h}$ denote the values of $w$ and $\pi_{t}$ respectively in HedgeWithQueries, while $w$ and $\pi_{t}$ refer to the variables in Algorithm 1. Then $w$ and $w^{h}$ evolve in the exact same way, so the distributions of $\pi_{t}$ and $\pi_{t}^{h}$ coincide. Thus by Lemma 5.4,
@@
$\operatorname*{\mathbb{E}}_{\mathbf{a}}\left[\sum_{t=1}^{T}\ell(t,\pi_{t})\right]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\ell(t,\tilde{\pi})\leq\frac{2\log|\tilde{\Pi}|}{p^{2}}=2T\varepsilon\log|\tilde{\Pi}|$
Since Lemma 5.4 holds for any loss function, the bound above holds for any $\mathbf{s}\in S^{T}$, so the bound also holds in expectation over $\mathbf{s}\sim\boldsymbol{\mathcal{D}}$ (which is needed for Case 1). Next, observe that $|V_{T}|=\sum_{t=1}^{T}\mathbf{1}(\pi_{t}(s_{t})\neq\pi^{m}(s_{t}))=\sum_{t=1}^{T}\ell(t,\pi_{t})$, so
$\operatorname*{\mathbb{E}}_{\mathbf{a}}[|V_{T}|]\leq 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))$
Case 1: Since $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of $\Pi$, we have
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))\right]\leq$ $\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\sup_{\pi\in\Pi}\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi(s_{t}))\right]$
$\in$ $\ O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$
with the first step due to $\pi^{m}\in\Pi$ and the second step due to Lemma 5.3. The last component we need is that $|\tilde{\Pi}|\leq(41/\varepsilon)^{d}$ by construction (and such a $\tilde{\Pi}$ is guaranteed to exist by Lemma 5.1). Combining the above inequalities and taking the expectation over $\mathbf{s}\sim\boldsymbol{\mathcal{D}}$ (in addition to the randomness of the algorithm), we get
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}},\,\mathbf{a}}\,[|V_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))\right]$
$\leq$ $\ 2dT\varepsilon\log(41/\varepsilon)+O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$
$\in$ $\ O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
Case 2: Since $\tilde{\Pi}$ is an adversarial cover of $\Pi$ and $\pi^{m}\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))=0$. Since $|\tilde{\Pi}|\leq(eT/d)^{d}$ (with such a $\tilde{\Pi}$ guaranteed to exist by Lemma 5.2),
$\operatorname*{\mathbb{E}}_{\mathbf{a}}[|V_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))$
$\leq$ $\ 2T\varepsilon d\ln(eT/d)$
$\in$ $\ O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
@@
Fix any $t\in[T]$. If $t\in Q_{T}$, then $\mu(s_{t},a_{t})=\mu^{m}(s_{t})$, so assume $t\not\in Q_{T}$. Let $(s^{\prime},a^{\prime})=\operatorname*{arg\,min}_{(s,a)\in X_{t}:\pi_{t}(s_{t})=a}||s_{t}-s||$. Since $t\not\in Q_{T}$, we must have $||s_{t}-s^{\prime}||\leq\varepsilon^{1/n}$.
We have $a^{\prime}=\pi^{m}(s^{\prime})$ by construction of $X_{t}$ and $\pi_{t}(s_{t})=a^{\prime}$ by construction of $a^{\prime}$. Combining these with the local generalization assumption, we get
$\mu(s_{t},a_{t})=\mu(s_{t},\pi_{t}(s_{t}))=\mu(s_{t},\pi^{m}(s^{\prime}))\geq\mu^{m}(s_{t})-L||s_{t}-s^{\prime}||\geq\mu^{m}(s_{t})-L\varepsilon^{1/n}$
@@
Assume $x_{1},\dots,x_{T},y_{1},\dots,y_{T}\in[0,1]$ and $x_{t}\geq y_{t}$ for all $t\in[T]$. Then
$\prod_{t=1}^{T}x_{t}-\prod_{t=1}^{T}y_{t}\leq\sum_{t=1}^{T}x_{t}-\sum_{t=1}^{T}y_{t}$
Proof.
We proceed by induction on $T$. The claim is trivially satisfied for $T=1$, so suppose $T>1$ and assume that $\prod_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}y_{t}\leq\sum_{t=1}^{T-1}x_{t}-\sum_{t=1}^{T-1}y_{t}$. Then
$\sum_{t=1}^{T}x_{t}-\sum_{t=1}^{T}y_{t}-\prod_{t=1}^{T}x_{t}+\prod_{t=1}^{T}y_{t}=$ $\ x_{T}\sum_{t=1}^{T-1}x_{t}-y_{T}\sum_{t=1}^{T-1}y_{t}-x_{T}\prod_{t=1}^{T-1}x_{t}+y_{T}\prod_{t=1}^{T-1}y_{t}$
$=$ $\ x_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\right)-y_{T}\left(\sum_{t=1}^{T-1}y_{t}-\prod_{t=1}^{T-1}y_{t}\right)$
Since $T>1$ and $x_{t}\in[0,1]$ for all $t\in[T]$, we have $\sum_{t=1}^{T-1}x_{t}\geq x_{1}\geq\sum_{t=1}^{T-1}x_{t}$. Thus $\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\geq 0$. Combining this with $x_{T}\geq y_{T}$, we get
$\sum_{t=1}^{T}x_{t}-\sum_{t=1}^{T}y_{t}-\prod_{t=1}^{T}x_{t}+\prod_{t=1}^{T}y_{t}=$ $\ x_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\right)-y_{T}\left(\sum_{t=1}^{T-1}y_{t}-\prod_{t=1}^{T-1}y_{t}\right)$
$\geq$ $\ y_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\right)-y_{T}\left(\sum_{t=1}^{T-1}y_{t}-\prod_{t=1}^{T-1}y_{t}\right)$
$=$ $\ y_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}-\sum_{t=1}^{T-1}y_{t}+\prod_{t=1}^{T-1}y_{t}\right)$
[…]
The last step is due to $y_{T}\geq 0$ and our assumption of $\prod_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}y_{t}\leq\sum_{t=1}^{T-1}x_{t}-\sum_{t=1}^{T-1}y_{t}$. ∎
Lemma B.4.
@@
$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
@@
We first claim that $a_{t}=\pi^{m}(s_{t})$ for all $t\not\in V_{T}$. If $t\in Q_{T}$, the claim is immediate; if not, we have $a_{t}=\pi_{t}(s_{t})$, and $\pi_{t}(s_{t})=\pi^{m}(s_{t})$ due to $t\not\in V_{T}$. Thus $\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))=\mu^{m}(s_{t})$ for $t\not\in V_{T}$.
We next claim that $\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\leq L\varepsilon^{1/n}$ for all $t\in[T]$. If $\mu(s_{t},a_{t})\leq\mu^{m}(s_{t})$, this follows from Lemma B.2. If $\mu(s_{t},a_{t})>\mu^{m}(s_{t})$, then $\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))=0\leq L\varepsilon^{1/n}$. Therefore
$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq$ $\ \sum_{t\in V_{T}}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
$\leq$ $\ \sum_{t\in V_{T}}L\varepsilon^{1/n}$
$=$ $\ |V_{T}|L\varepsilon^{1/n}$
Now let $x_{t}=\mu^{m}(s_{t})$ and $y_{t}=\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
$\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
Since $\mu(s_{t},a_{t})\geq\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$, we have
$R_{T}=$ $\ \prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
$\leq$ $\ |V_{T}|L\varepsilon^{1/n}$
Taking the expectation and applying Lemma B.1 completes the proof. ∎
@@
Let $(K,||\cdot||)$ be a normed vector space and let $\delta>0$. Then $W\subseteq K$ is a $\delta$-packing of $K$ if for all $x,y\in W$, $||x-y||>\delta$. The $\delta$-packing number of $K$, denoted $\mathcal{M}(K,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $K$.
In this paper, we only consider the Euclidean distance norm, so we just write $M(K,||\cdot||,\delta)=M(K,\delta)$.
@@
If $K\subset\mathbb{R}^{n}$ is convex, bounded, and contains a ball with radius $\delta>0$, then
$\mathcal{M}(K,\delta)\leq\frac{3^{n}\operatorname*{vol}(K)}{\delta^{n}\operatorname*{vol}(B)}$
where $B$ is a unit ball.
Lemma B.6 (Jung’s Theorem [22]).
[…]